OpenStudy (anonymous):

Rationalize the denominator of sqrt -49 over (7 - 2i) - (4 + 9i)

OpenStudy (anonymous):

I know that the sqrt of -49 factors into i49.

OpenStudy (anonymous):

i sqrt 49 is what I meant

OpenStudy (jdoe0001):

\(\bf \cfrac{\sqrt{-49}}{(7 - 2i) - (4 + 9i)}\\ \quad \\ \textit{notice that }\quad \sqrt{-49}\implies \sqrt{49}\times \sqrt{-1}\\ \quad \\ \textit{let us use the }conjugate\textit{ of the denominator } (7 - 2i) + (4 + 9i)\\ \cfrac{\sqrt{-49}}{(7 - 2i) - (4 + 9i)} \times \cfrac{(7 - 2i) + (4 + 9i)}{(7 - 2i) + (4 + 9i)}\\ \quad \\ \cfrac{\sqrt{-49}[(7 - 2i) + (4 + 9i)]}{[(7 - 2i) - (4 + 9i)]\quad [(7 - 2i) + (4 + 9i)]} \\ \quad \\ \textit{keep in mind that } (a-b)(a+b) = (a^2-b^2)\\ \quad \\ \cfrac{\sqrt{-49}[(7 - 2i) + (4 + 9i)]}{[(7 - 2i) - (4 + 9i)]\quad [(7 - 2i) + (4 + 9i)]}\\\quad \\ \implies \cfrac{\sqrt{-49}[(7 - 2i) + (4 + 9i)]}{(7 - 2i)^2 - (4 + 9i)^2}\)

OpenStudy (anonymous):

It kind of makes sense, but I thought you had to make sqrt -49 into a positive

OpenStudy (jdoe0001):

you'd eventually, yes

OpenStudy (anonymous):

I'm still confused though :l

OpenStudy (anonymous):

oh okay

OpenStudy (jdoe0001):

ok.... where? you still need to simplify it some, expand the bottom and top

OpenStudy (jdoe0001):

.... ohhh... smokes... wait a second

OpenStudy (anonymous):

I don't understand the 3rd step. I just don't remember all of that in my lessons

OpenStudy (anonymous):

Okay (:

OpenStudy (jdoe0001):

... all that was so superfluos.... shoot the denominator is easier simplifyable

OpenStudy (jdoe0001):

lemme redo all that hehe

OpenStudy (anonymous):

lol Okayy

OpenStudy (jdoe0001):

\(\bf \cfrac{\sqrt{-49}}{(7 - 2i) - (4 + 9i)}\\ \quad \\ \textit{notice that }\quad \sqrt{-49}\implies \sqrt{49}i\implies \sqrt{7^2}i\implies 7i\\ \quad \\ \cfrac{\sqrt{-49}}{7 - 2i - 4 - 9i}\implies \cfrac{7i}{3 - 7i}\\ \quad \\ \textit{now let's use the conjugate of the denominator of }\quad 3+7i\\ \quad \\ \cfrac{7i}{3 - 7i}\times \cfrac{3+7i}{3+7i} \implies \cfrac{7i(3+7i)}{(3-7i)(3+7i)}\\ \quad \\ \textit{now recall that }\quad (a-b)(a+b) = (a^2-b^2)\\ \quad \\ \cfrac{7i(3+7i)}{(3-7i)(3+7i)}\implies \cfrac{7i(3+7i)}{(3)^2-(7i)^2}\implies \cfrac{21i+49i^2}{9-7^2i^2}\)

OpenStudy (jdoe0001):

acck... hold.... I have a slight error there :( -2i -9i = -11i lemme fix that quick

OpenStudy (anonymous):

its okay (: I forgot about seeing if -49 had any factors

OpenStudy (jdoe0001):

\(\bf \cfrac{\sqrt{-49}}{(7 - 2i) - (4 + 9i)}\\ \quad \\ \textit{notice that }\quad \sqrt{-49}\implies \sqrt{49}i\implies \sqrt{7^2}i\implies 7i\\ \quad \\ \cfrac{\sqrt{-49}}{7 - 2i - 4 - 9i}\implies \cfrac{7i}{3 - 11i}\\ \quad \\ \textit{now let's use the conjugate of the denominator of }\quad 3+11i\\ \quad \\ \cfrac{7i}{3 - 11i}\times \cfrac{3+11i}{3+11i} \implies \cfrac{7i(3+11i)}{(3-11i)(3+11i)}\\ \quad \\ \textit{now recall that }\quad (a-b)(a+b) = (a^2-b^2)\\ \quad \\ \cfrac{7i(3+11i)}{(3-11i)(3+11i)}\implies \cfrac{7i(3+11i)}{(3)^2-(11i)^2}\implies \cfrac{21i+77i^2}{9-11^2i^2}\)

OpenStudy (anonymous):

That's the final answer?

OpenStudy (jdoe0001):

and when expanding the denominator, don't forget that \(\bf i^2\implies \sqrt{-1}\times \sqrt{-1}\implies \sqrt{(-1)^2}\implies -1\)

OpenStudy (jdoe0001):

\(\bf \cfrac{21i+77i^2}{9-11^2i^2}\implies \cfrac{21i+77(-1)}{9-11^2(-1)}\)

OpenStudy (anonymous):

Okay, but that's not any of the answer choices

OpenStudy (jdoe0001):

well, what would you get for \(\bf \cfrac{21i+77(-1)}{9-11^2(-1)}\)

OpenStudy (anonymous):

what do you mean?

OpenStudy (jdoe0001):

if you expand both top and bottom

OpenStudy (jdoe0001):

what's \(\bf 77 \times -1\quad ?\) what's \(\bf -11^2 \times -1\quad ?\)

OpenStudy (anonymous):

Just making sure, is x suppose to be a variable or a multiplication sign? lol

OpenStudy (jdoe0001):

multiplication :) or maybe better \(\bf 77 \cdot -1\\ -11^2 \cdot -1 \quad ?\)

OpenStudy (anonymous):

That's better lol -77 and 11^2?

OpenStudy (jdoe0001):

right and 11^2 = 121 so we can say that \(\bf \cfrac{21i+77i^2}{9-11^2i^2}\implies \cfrac{21i+77(-1)}{9-11^2(-1)} \implies \cfrac{21i-77}{9+121} \implies \cfrac{-77+21i}{130} \)

OpenStudy (anonymous):

Oh, okay! That makes sense!

OpenStudy (anonymous):

Can you help me out with another one? It's simple. I just forgot the first step lol

OpenStudy (jdoe0001):

ok

OpenStudy (anonymous):

Simplify 4 sqrt 192

OpenStudy (anonymous):

I have to factor 192 right?

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

okay, that's all I needed to know. Thanks so much!

OpenStudy (jdoe0001):

 \(\bf 192 \implies 2\cdot 2\cdot 2\cdot 2\cdot 2\cdot 2\cdot 3\implies 2^2\cdot 2^2\cdot 2^2\cdot 3\\\implies (2^2)^3\cdot 3\implies (2^3)^2\cdot 3\\ \quad \\ 4\sqrt{192} \implies 4\sqrt{(2^3)^2\cdot 3}\)